Hi My little cousin needs help with this homework problem, I tried helping her but I can't seem to get the right answer.
Three ducks and two ducklings weigh 77 kg Four ducks and three ducklings weigh 107 kg All ducks weigh the same and all ducklings weigh the same What is the weight of two ducks and one duckling?
This is how I set up the problem
D=Duck
d=Duckling
(3D + 2d = 77)
(4D + 3d = 107)

Answer by checkley77(12844)

D=Duck
d=Duckling
(3D + 2d = 77)
3D=77-2d
D=(77-2d)/3
-----------------
Replace D with (77-2d)/3 & solve for d.
(4D + 3d = 107)
4(77-2d)/3+3d=107
(308-8d)/3+3d=107
(308-8d+3*3d)/3=107
(308-8d+9d)/3=107
(308+d)/3=107 cross multiply.
308+d=3*107
308+d=321
d=321-308
d=13 ans.
D=(77-2*13)/3
D=(77-26)/3
D=51/3
D=17 ans.
Proof:
4*17+3*13=107
68+39=107
107=107